# Dynamic and Distributed Routing in IoT Networks based on Multi-Objective Q-Learning

### v1

@@ V Distributed DPQ-Learning
VI Experimental Results
@@
    Reinforcement-Learning-Based Energy Efficient Control and Routing Protocol for Wireless Sensor Networks (RLBEEP) [13]
@@
    Reinforcement Learning for Life Time Optimization (R2LTO) [14]
@@
    Reinforcement Learning Based Routing (RLBR) [15]
@@
VI-A The Simulation Setup and Parameters
@@
The variation of $\epsilon$-values and preference variation are shown in Figure 2. Since experiments 1 and 2 are characterized by sequential exploration and exploitation, the $\epsilon$ value (Figure 2(a)) in the proposed algorithm is $1$ in the first $1000$ episodes (indicating full exploration and no exploitation) and 00 afterward (indicating no exploration and only exploitation). In experiments $3$ and $4$ (Figure 2(b)), the $\epsilon$ is set as $1$ in the beginning and then linearly reduced to 00, thus achieving the desired simultaneous exploration and exploitation. In all the experiments, the $\epsilon$ values for the SMORLR baseline are reset to $1$ each time the decision-maker changes the preference. Then, it linearly reduces to 00, encouraging fresh exploration after each preference-change. The frequency of preference variation is a thousand episodes in experiments $1$ and $3$ (Figure 2(c)), and each episode in experiments $2$ and $4$ (Figure 2(d)). The two different preference variation patterns are considered to study the effect of the frequency of preference changes on performance.
@@
We compare the performance of the proposed Distributed DPQ-Learning routing algorithm with the six baselines on three vital metrics: overall reward, packet delivery, and energy consumption. The following subsections provide detailed analysis of these three metrics under sequential and simultaneous exploration-exploitation scenarios.
VI-B Sequential Exploration and Exploitation
@@ VI-B1 Overall Reward
Figure 3 shows the performance comparison of $r^{\texttt{Overall}}(s,a)$ in the sequential exploration-exploitation scheme between the proposed and existing works. Figure 3(a) and Figure 3(b) shows experiment 1’s episodic and accumulated rewards, respectively. Figure 3(a) shows that the proposed Distributed DPQ-Learning routing algorithm immediately adapts according to the preference and gives high performance even if the preferences change. This superior performance is attributed to the Pareto front, which is learned during the exploration phase. In contrast SMORLR starts learning the policy suited to the new preference after each preference shift. Although it focuses on the same objectives, it still does not learn enough to outperform the proposed dynamic protocol after 1000 episodes of learning. Static protocols perform poorly in environments where preferences change dynamically. Algorithms like R2LTO, RLBR, and RLBEEP, which are primarily energy-focused and static, display performance that is relatively similar to one another for most preferences. However, they underperform by an average of 102% compared to the proposed approach. During the intervals between $1000-2000$ and $6000-7000$ episodes, these algorithms begin to show improvement, as the preference weight, $\beta$, shifts to $0.1$ and $0.2$, giving higher priority to energy efficiency–an objective for which they are optimized. RSPAR underperforms by 263% on average compared to the proposed method, while FMOLD, despite being a multi-objective approach, is 266% less efficient. Moreover, they don’t seem to be learning efficient policies with time. However, the proposed algorithm performs better immediately (from the beginning of these intervals) because it utilizes the Pareto front learned during exploration.
@@
The cumulative reward is shown in Figure 3(b), and confirm superiority in the proposed distributed DPQ-Learning algorithm over baseline approaches over the long run. Through the episodes, this outperformance becomes increasingly evident. Finally, after 10000 episodes, we can notice that the accumulated return for the proposed method is better than SMORLR by 2.3×2.3$\times$2.3 ×, RLBEEP by 5.12×5.12$\times$5.12 ×, R2LTO and FMOLD by 5.25×5.25$\times$5.25 ×, RLBR by 6×6$\times$6 ×, and RSPAR by 16.8×16.8$\times$16.8 ×. Similarly, in experiment 2, preferences were randomly changed in each episode and the measured performance is shown in Figure 3(c). Here we noticed that the accumulated returns of the proposed routing protocol are consistently higher than baselines, and the difference becomes larger with more episodes. At the end of $100$ episodes, the accumulated return of the proposed algorithm is around $43$ which is $\times3$ of RSPAR, 4.3×4.3$\times$4.3 × of SMORLR, 6.14×6.14$\times$6.14 × of RLBR, 7.16×7.16$\times$7.16 × of R2LTO, 8.6×8.6$\times$8.6 × of FMOLD and 43×43$\times$43 × of RLBEEP. Our observations confirm the proposed method consistently outperforms all baselines and adapts to this efficient performance irrespective of preference changes.
VI-B2 Energy
@@
Figure 4(a) compares episodic energy consumption (in mJ) across baseline algorithms over $10,000$ episodes, showing their performance trends. The Y-axis in Figure 4(a) is logarithmic, ranging from 00 to $1$ mJ. The proposed distributed DPQ-Learning routing protocol leads to minimal episodic energy consumption whenever a higher preference is given to the energy objective while performing reasonably well even when the preference for the energy objective is low. On average, the proposed distributed DPQ-Learning routing protocol shows about $80-90$% lower energy consumption than baseline algorithms. SMORLR consumes around $150-200\%$ more energy than the proposed algorithm over time. However, with each preference change, it starts to learn to converge closer to the proposed algorithm gradually. However, the more often preferences change, the inefficient this approach becomes. R2LTO and RLBR tend to stabilize close to $0.3$ mJ of episodic energy consumption, with about $50\%$ higher consumption than the proposed method. RLBEEP spikes at first and then stabilizes around $0.4$ mJ. It has approximately $100\%$ higher consumption than the proposed algorithm. FMOLD stabilizes near $0.6$ mJ, about $3$ times the energy of the distributed DPQ-Learning approach. The energy consumption is the highest among the four, but the consistency is better than RSPAR. The lowest energy efficiency can be seen in [RSPAR], which spikes above $1$ mJ initially and stabilizes around $0.8$ mJ. Its energy consumption is 4×4$\times$4 × to 5×5$\times$5 × higher than the proposed algorithm. Although RLBEEP, R2LTO, and RLBR learn to perform well or better than the proposed method for certain sets of preferences, but their overall performance over a long period is lesser than the proposed method.
@@
The cumulative effects over the long run can be more clearly seen in Figure 4(b), which depicts the cumulative energy consumption over episodes. The figure shows that the proposed algorithm is the most energy-efficient, consuming $30-50\%$ lesser energy than SMORLR, RLBEEP, RLBR, and R2LTO, and around 4×4$\times$4 × and 3.4×3.4$\times$3.4 × less than RSPAR and FMOLD respectively by episode $10000$. While SMORLR, RLBEEP, R2LTO, and RLBR perform moderately well, RSPAR and FMOLD exhibit the highest cumulative energy consumption, highlighting their inefficiency. However, the proposed routing algorithm still consumes less energy than any baseline algorithm. Similarly we continue our evaluations in Experiment 2, where preferences were randomly changed in each episode. As shown in Figure 4(c), the cumulative energy consumption for the proposed routing protocol is lesser than those of the baseline methods, with the gap widening as more episodes progress. By the end of 100 episodes, we find that the cumulative energy consumption of RSPAR and RLBEEP is 3.87×3.87$\times$3.87 ×, FMOLD is 3.53×3.53$\times$3.53 ×, RLBR is 2×2$\times$2 ×, R2LTO is 1.73×1.73$\times$1.73 ×, and SMORLR is 1.67×1.67$\times$1.67 × that of the proposed routing scheme. Therefore, regardless of the exploration policy or preference variation patterns, the proposed method leads to a substantially lesser amount of energy consumption than the baseline algorithms.
VI-B3 Packet Delivery
@@
Figure 5(a) and Figure 5(b) depict the variation in episodic and cumulative packet deliveries, respectively. Figure 5(a) shows that the proposed distributed DPQ-Learning routing algorithm outperforms all the baselines for episodes ($0-1000,2000-4000,5000-6000,7000-10000$) where the preference for the PDR objective is more than 0.5. RSPAR and FMOLD tend to hover around episodic PDR of $0.4-0.5$ and $0.3-0.4$ respectively. SMORLR tends to learn with time to maximize the PDR for each new preference vector. However, its peak performances are $50\%$ lower than the proposed method’s consistent performances. RLBEEP, R2LTO, and RLBR focus on the energy aspect of the system and did not consider PDR. As a result, they consistently ignore this objective and performing worse. At the beginning, R2LTO and RLBR seem to perform fairly well against PDR, but eventually, they start performing poorly as well.
@@
(b) Experiment 1: Comparison of no. of packets delivered
(c) Experiment 2: Total Packets Delivered v/s Episodes
@@
We analyze cumulative packet delivery over a long run as shown in Figure 5(b), and the proposed method consistently outperforms all the baselines except for some initial fluctuations durations. As time passes, the difference between performances becomes more evident. At the end of $10000$ episodes, the proposed method has better performance than RSPAR by around 1.78×1.78$\times$1.78 ×, FMOLD and RLBEEP by 3×3$\times$3 ×, and R2LTO and RLBR by around 3.5×3.5$\times$3.5 ×. Similar results were observed in Experiment 2, where preferences were randomly changed in each episode as shown in Figure 5(c). We observe here that the total number of packets delivered by the proposed work consistently exceeds those delivered by the baseline method, and that the gap widens as more episodes are completed. By the end of $100$ episodes, the proposed algorithm’s accumulated return reaches approximately $68$, which is 1.41×1.41$\times$1.41 × that of RSPAR, 2.6×2.6$\times$2.6 × greater than FMOLD, 2.62×2.62$\times$2.62 × better than R2LTO and RLBR, 2.83×2.83$\times$2.83 × higher than SMORLR, and 3.1×3.1$\times$3.1 × higher than RLBEEP. Consequently, regardless of the exploration policy or preference variation patterns, the proposed method outperforms the baseline algorithms in terms of PDR.
VI-C Simultaneous Exploration and Exploitation

### v2

@@ V Distributed DPQ-Learning
VI Performance Guarantees for the GIP
In this section, we quantify the approximation accuracy of the GIP policy. As discussed in Section IV, the learning procedure converges, at the grid points $\beta\in\mathcal{B}$, to the corresponding optimal $Q$-tables $Q_{\beta}$. For any intermediate $\beta\notin\mathcal{B}$, Algorithm 2 constructs an interpolated $Q$-table from its two neighboring grid points and acts greedily with respect to it. We establish a uniform approximation bound for this policy with respect to the optimal policy at the same $\beta$. The guarantee is controlled solely by standard problem constants and the grid resolution; thus, refining the grid and improving the per-grid estimates makes the GIP policy uniformly near-optimal over the entire $\beta$-range.
Our bounds depend on two problem-level parameters. First, the worst-case expected episode length under optimal control, captured by a single constant that applies to all $\beta$ and all initial states. Let $K$ denote the (random) episode length, i.e., the number of time steps from an initial state $s$ until termination. For each $\beta$, define the worst $\beta$-optimal expected episode length from state $s$ by
$H_{\beta}(s)\;\triangleq\;\sup_{\pi\in\Pi_{\beta}^{*}}\,\mathbb{E}_{s}^{\pi}[K],\qquad H\;\triangleq\;\sup_{\beta\in\mathcal{B}}\;\sup_{s\in\mathcal{S}}\,H_{\beta}(s).$
We take the supremum over optimal policies $\Pi_{\beta}^{*}$ so that the bound is independent of tie-breaking among optimal actions. Under standard stochastic shortest-path assumptions (proper optimal policies and bounded rewards), $H<\infty$. Second, our bounds will depend on the difference between the two objectives, or rewards. In particular, define
$\Gamma\triangleq\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}|r^{\rm{Energy}}(s,a)-r^{\texttt{PDR}}(s,a)|.$ (6)
We now prepare to state the approximation bound for the GIP policy. We analyze the approximation error after running Algorithm 1 and assume that the $Q$-tables learned at the grid points are accurate to within a common tolerance $\epsilon$, formalized in the following assumption.
Assumption 1.
For each $\beta\in\mathcal{B}$ we have an estimate $\widehat{Q}_{\beta}$ satisfying
$\|\widehat{Q}_{\beta}-Q_{\beta}\|_{\infty}\leq\varepsilon,$
where $Q_{\beta}$ is the optimal Q-table for preference indicator $\beta$.
Then the GIP policy for any $\beta\in[0,1]$ is a greedy policy with respect an interpolation of $\hat{Q}_{\underline{\beta}}$ and $\hat{Q}_{\overline{\beta}}$ where $\underline{\beta}\triangleq\max\{b\in\mathcal{B}|b\leq\beta\}$ and $\overline{\beta}\triangleq\min\{b\in\mathcal{B}|b\geq\beta\}$. In particular, it is greedy with respect to the Q-table
$Q_{\beta}^{\texttt{Int}}\triangleq\rho\hat{Q}_{\underline{\beta}}+(1-\rho)\hat{Q}_{\overline{\beta}},$ (7)
where $\rho\triangleq(b-\underline{\beta})/(\overline{\beta}-\underline{\beta})$. Meaning that
$\pi_{\beta}^{\texttt{Greedy}}(s):=\underset{a\in\mathcal{A}}{\text{argmax}}~Q_{\beta}^{\texttt{Int}}(s,a).$
Theorem 1.
Under Assumption 1, for any $\beta\in[0,1]$ the interpolation policy in Eq. 7
$Q_{\beta}^{\texttt{Int}}-Q_{\beta}||_{\infty}\leq\epsilon+\Gamma(H+1)(\overline{\beta}-\underline{\beta})$
where $\underline{\beta}\triangleq\max\{b\in\mathcal{B}|b\leq\beta\}$ and $\overline{\beta}\triangleq\min\{b\in\mathcal{B}|b\geq\beta\}$.
What this bound means is: if the $Q$-learning in Algorithm 1 runs at the grid points and produces approximate estimates of the optimal $Q$-tables, which is ensured under standard conditions (e.g., sufficient exploration and appropriate stepsizes), then for any preference $\beta$ the policy that is greedy with respect to the interpolated table is guaranteed to be near–optimal. In particular, its deviation from the optimal $Q_{\beta}$ is uniformly controlled across all $\beta\in[0,1]$ by the sum of the per–grid learning accuracy $\epsilon$ and a term that depends only on the local grid gap $(\overline{\beta}-\underline{\beta})$ and fixed problem constants. Practically, this provides a simple design rule: choose the grid resolution to meet a target accuracy, trading off computational cost (number of grid points) against uniform suboptimality of the GIP policy.
VI-A Proof of Theorem 1
To prove Theorem 1, we first establish a basic regularity property: the mapping $\beta\mapsto Q_{\beta}$ is Lipschitz continuous.
Lemma 1.
The optimal $Q$-table varies at most linearly with $\beta$:
$\|Q_{\beta}-Q_{\beta^{\prime}}\|_{\infty}\;\leq\;\Gamma\,(H+1)\,|\beta-\beta^{\prime}|\qquad\text{for all }\beta,\beta^{\prime}\in[0,1],$
where $\Gamma$ is the reward-variation constant and $H$ is the worst-case $\beta$-optimal expected episode length.
Proof.
Recall that the scalarized reward for $\beta\in[0,1]$ is a convex combination
$r_{\beta}(s,a)=\beta r^{\rm{Energy}}(s,a)+(1-\beta)r^{\texttt{PDR}}(s,a).$
Then, recalling the definition of $\Gamma$ in Eq. (6), for all $(s,a)$ and all $\beta,\beta^{\prime}\in[0,1]$,
$|r_{\beta}(s,a)-r_{\beta^{\prime}}(s,a)|=$ $|\beta-\beta^{\prime}|\;|r^{\rm{Energy}}(s,a)-r^{\texttt{PDR}}(s,a)|$
$\leq$ $\Gamma\,|\beta-\beta^{\prime}|.$
Fix a stationary policy $\pi$ and an initial state $s$. Let $K$ denote (random) episode length until termination. We have that
$\big|V_{\beta}^{\pi}(s)-V_{\beta^{\prime}}^{\pi}(s)\big|=$ $\Big|\mathbb{E}_{s}^{\pi}\Big[\sum_{t=0}^{K-1}\big(r_{\beta}(s_{t},a_{t})-r_{\beta^{\prime}}(s_{t},a_{t})\big)\Big]\Big|$
$\leq$ $\Gamma\,|\beta-\beta^{\prime}|\,\mathbb{E}_{s}^{\pi}[K].$
Let $\pi_{\beta}\in\Pi_{\beta}^{*}$ be $\beta$-optimal. Then, for any $s$,
$V_{\beta}(s)-V_{\beta^{\prime}}(s)\leq$ $V_{\beta}^{\pi_{\beta}}(s)-V_{\beta^{\prime}}^{\pi_{\beta}}(s)\leq\Gamma\,|\beta-\beta^{\prime}|\,\mathbb{E}_{s}^{\pi_{\beta}}[K]$
$\leq$ $\Gamma\,H\,|\beta-\beta^{\prime}|.$
Swapping $\beta$ and $\beta^{\prime}$ gives the reverse inequality; hence
$\|V_{\beta}-V_{\beta^{\prime}}\|_{\infty}\leq\Gamma\,H\,|\beta-\beta^{\prime}|.$
Finally, for any $(s,a)$ and random next state $s^{\prime}\sim T(s,a)$
$|Q_{\beta}(s,a)-Q_{\beta^{\prime}}(s,a)|$ $\leq\big|r_{\beta}(s,a)-r_{\beta^{\prime}}(s,a)\big|$
$~~~~~~~~~~+\Big|\mathbb{E}\big[V_{\beta}(s^{\prime})-V_{\beta^{\prime}}(s^{\prime})\mid s,a\big]\Big|$
$\leq\Gamma\,|\beta-\beta^{\prime}|+\|V_{\beta}-V_{\beta^{\prime}}\|_{\infty}$
$\leq\Gamma\,(1+H)\,|\beta-\beta^{\prime}|.$
Taking the maximum over $(s,a)$ yields
$\|Q_{\beta}^{*}-Q_{\beta^{\prime}}^{*}\|_{\infty}\leq\Gamma\,(1+H)\,|\beta-\beta^{\prime}|.$
We are now ready to prove Theorem 1. By the triangle inequality, we can slit the error in the following two terms:
[…]
$~\;+\;\underbrace{\big\|\rho\,Q_{\underline{\beta}}+(1-\rho)\,Q_{\overline{\beta}}-Q_{\beta}\big\|_{\infty}}_{T_{2}}.$
To bound $T_{1}$ we use triangle inequality and Assumption 1 to obtain $T_{1}\;\leq\;\rho\,\varepsilon+(1-\rho)\,\varepsilon\;=\;\varepsilon.$ To bound $T_{2}$, we note that by Lemma 1, $\|Q_{\underline{\beta}}-Q_{\beta}\|_{\infty}\leq\Gamma(H+1)\,|\beta-\underline{\beta}|$ and $\|Q_{\overline{\beta}}-Q_{\beta}\|_{\infty}\leq\Gamma(H+1)\,|\overline{\beta}-\beta|$. Hence,
$T_{2}$ $\;\leq\;\rho\,\Gamma(H+1)\,|\beta-\underline{\beta}|+(1-\rho)\,\Gamma(H+1)\,|\overline{\beta}-\beta|$
$\;\leq\;\Gamma(H+1)\,(\overline{\beta}-\underline{\beta}).$
Combining these results, yields the result $Q_{\beta}^{\texttt{Int}}-Q_{\beta}||_{\infty}\leq T_{1}+T_{2}\leq\epsilon+\Gamma(H+1)\,(\overline{\beta}-\underline{\beta})$
VII Experimental Results
@@
    Reinforcement-Learning-Based Energy Efficient Control and Routing Protocol for Wireless Sensor Networks (RLBEEP) [abadi2022rlbeep]
@@
    Reinforcement Learning for Life Time Optimization (R2LTO) [bouzid2020efficient]
@@
    Reinforcement Learning Based Routing (RLBR) [guo2019optimizing]
@@
VII-A The Simulation Setup and Parameters
@@
The variation of $\epsilon$-values and preference variation are shown in Figure 2. Since experiments 1 and 2 are characterized by sequential exploration and exploitation, the $\epsilon$ value (Figure 2(a)) in the proposed algorithm is $1$ in the first $1000$ episodes (indicating full exploration and no exploitation) and 0 afterward (indicating no exploration and only exploitation). In experiments $3$ and $4$ (Figure 2(b)), the $\epsilon$ is set as $1$ in the beginning and then linearly reduced to 0, thus achieving the desired simultaneous exploration and exploitation. In all the experiments, the $\epsilon$ values for the SMORLR baseline are reset to $1$ each time the decision-maker changes the preference. Then, it linearly reduces to 0, encouraging fresh exploration after each preference-change. The frequency of preference variation is a thousand episodes in experiments $1$ and $3$ (Figure 2(c)), and each episode in experiments $2$ and $4$ (Figure 2(d)). The two different preference variation patterns are considered to study the effect of the frequency of preference changes on performance.
@@
In the simulations, transmission priorities are represented by the preference parameter $\beta$, which adjusts the balance between packet delivery ratio (PDR) and energy efficiency. Lower $\beta$ values (e.g., $\beta=0.1$) correspond to high-priority transmissions that favor higher PDR and reliability, even at the cost of increased energy consumption. In contrast, higher $\beta$ values (e.g., $\beta=0.9$) represent low-priority transmissions that prioritize energy savings over delivery reliability. We compare the performance of the proposed Distributed DPQ-Learning routing algorithm with the six baselines on three vital metrics: overall reward, packet delivery, and energy consumption. The following subsections provide detailed analysis of these three metrics under sequential and simultaneous exploration-exploitation scenarios.
VII-B Sequential Exploration and Exploitation
@@ VII-B1 Overall Reward
Figure 3 shows the performance comparison of $r^{\texttt{Overall}}(s,a)$ in the sequential exploration-exploitation scheme between the proposed and existing works. Figure 3(a) and Figure 3(b) shows experiment 1’s episodic and accumulated rewards, respectively. Figure 3(a) shows that the proposed Distributed DPQ-Learning routing algorithm immediately adapts according to the preference and gives high performance even if the preferences change. This superior performance is attributed to the Pareto front, which is learned during the exploration phase. In contrast SMORLR starts learning the policy suited to the new preference after each preference shift. Although it focuses on the same objectives, it still does not learn enough to outperform the proposed dynamic protocol after 1000 episodes of learning. Static protocols perform poorly in environments where preferences change dynamically. Algorithms like R2LTO, RLBR, and RLBEEP, which are primarily energy-focused and static, display performance that is relatively similar to one another for most preferences. However, they underperform by an average of 102% compared to the proposed approach. During the intervals between 100020001000-2000 and 600070006000-7000 episodes, these algorithms begin to show improvement, as the preference weight, $\beta$, shifts to $0.1$ and $0.2$, giving higher priority to energy efficiency–an objective for which they are optimized. RSPAR underperforms by 263% on average compared to the proposed method, while FMOLD, despite being a multi-objective approach, is 266% less efficient. Moreover, they don’t seem to be learning efficient policies with time. However, the proposed algorithm performs better immediately (from the beginning of these intervals) because it utilizes the Pareto front learned during exploration.
@@
The cumulative reward is shown in Figure 3(b), and confirm superiority in the proposed distributed DPQ-Learning algorithm over baseline approaches over the long run. Through the episodes, this outperformance becomes increasingly evident. Finally, after 10000 episodes, we can notice that the accumulated return for the proposed method is better than SMORLR by $2.3\times$, RLBEEP by $5.12\times$, R2LTO and FMOLD by $5.25\times$, RLBR by $6\times$, and RSPAR by $16.8\times$. Similarly, in experiment 2, preferences were randomly changed in each episode and the measured performance is shown in Figure 3(c). Here we noticed that the accumulated returns of the proposed routing protocol are consistently higher than baselines, and the difference becomes larger with more episodes. At the end of $100$ episodes, the accumulated return of the proposed algorithm is around $43$ which is $3.3\times$ of RSPAR, $4.3\times$ of SMORLR, $6.14\times$ of RLBR, $7.16\times$ of R2LTO, $8.6\times$ of FMOLD and $43\times$ of RLBEEP. Our observations confirm the proposed method consistently outperforms all baselines and adapts to this efficient performance irrespective of preference changes.
VII-B2 Energy
@@
Figure 4(a) compares episodic energy consumption (in mJ) across baseline algorithms over $10,000$ episodes, showing their performance trends. The Y-axis in Figure 4(a) is logarithmic, ranging from 0 to $1$ mJ. The proposed distributed DPQ-Learning routing protocol leads to minimal episodic energy consumption whenever a higher preference is given to the energy objective while performing reasonably well even when the preference for the energy objective is low. On average, the proposed distributed DPQ-Learning routing protocol shows about 809080-90% lower energy consumption than baseline algorithms. SMORLR consumes around $150-200\%$ more energy than the proposed algorithm over time. However, with each preference change, it starts to learn to converge closer to the proposed algorithm gradually. However, the more often preferences change, the inefficient this approach becomes. R2LTO and RLBR tend to stabilize close to $0.3$ mJ of episodic energy consumption, with about $50\%$ higher consumption than the proposed method. RLBEEP spikes at first and then stabilizes around $0.4$ mJ. It has approximately $100\%$ higher consumption than the proposed algorithm. FMOLD stabilizes near $0.6$ mJ, about $3$ times the energy of the distributed DPQ-Learning approach. The energy consumption is the highest among the four, but the consistency is better than RSPAR. The lowest energy efficiency can be seen in RSPAR, which spikes above $1$ mJ initially and stabilizes around $0.8$ mJ. Its energy consumption is $4\times$ to $5\times$ higher than the proposed algorithm. Although RLBEEP, R2LTO, and RLBR learn to perform well or better than the proposed method for certain sets of preferences, but their overall performance over a long period is lesser than the proposed method.
@@
The cumulative effects over the long run can be more clearly seen in Figure 4(b), which depicts the cumulative energy consumption over episodes. The figure shows that the proposed algorithm is the most energy-efficient, consuming around $4\times$ and $3.4\times$ less than RSPAR and FMOLD respectively, $30-50\%$ lesser energy than SMORLR and RLBEEP, and $3-5\%$ lesser than RLBR, and R2LTO by episode $10000$. While SMORLR, RLBEEP, R2LTO, and RLBR perform moderately well, RSPAR and FMOLD exhibit the highest cumulative energy consumption, highlighting their inefficiency. However, the proposed routing algorithm still consumes less energy than any baseline algorithm. Similarly we continue our evaluations in Experiment 2, where preferences were randomly changed in each episode. As shown in Figure 4(c), the cumulative energy consumption for the proposed routing protocol is lesser than those of the baseline methods, with the gap widening as more episodes progress. By the end of 100 episodes, we find that the cumulative energy consumption of RSPAR and RLBEEP is $3.87\times$, FMOLD is $3.53\times$, RLBR is $2\times$, R2LTO is $1.73\times$, and SMORLR is $1.67\times$ that of the proposed routing scheme. Therefore, regardless of the exploration policy or preference variation patterns, the proposed method leads to a substantially lesser amount of energy consumption than the baseline algorithms.
VII-B3 Packet Delivery
@@
Figure 5(a) and Figure 5(b) depict the variation in episodic and cumulative packet deliveries, respectively. Figure 5(a) shows that the proposed distributed DPQ-Learning routing algorithm outperforms all the baselines for episodes (01000,20004000,50006000,7000100000-1000,2000-4000,5000-6000,7000-10000) where the preference for the PDR objective is more than 0.5. RSPAR and FMOLD tend to hover around episodic PDR of 0.40.50.4-0.5 and 0.30.40.3-0.4 respectively. SMORLR tends to learn with time to maximize the PDR for each new preference vector. However, its peak performances are $50\%$ lower than the proposed method’s consistent performances. RLBEEP, R2LTO, and RLBR focus on the energy aspect of the system and did not consider PDR. As a result, they consistently ignore this objective and performing worse. At the beginning, R2LTO and RLBR seem to perform fairly well against PDR, but eventually, they start performing poorly as well.
@@
(b) Exp. 1: Packets Delivered v/s Episodes
(c) Exp. 2: Packets Delivered v/s Episodes
@@
We analyze cumulative packet delivery over a long run as shown in Figure 5(b), and the proposed method consistently outperforms all the baselines except for some initial fluctuations durations. As time passes, the difference between performances becomes more evident. At the end of $10000$ episodes, the proposed method has better performance than RSPAR by around $1.78\times$, FMOLD and RLBEEP by $3\times$, and R2LTO and RLBR by around $3.5\times$. Similar results were observed in Experiment 2, where preferences were randomly changed in each episode as shown in Figure 5(c). We observe here that the total number of packets delivered by the proposed work consistently exceeds those delivered by the baseline method, and that the gap widens as more episodes are completed. By the end of $100$ episodes, the proposed algorithm’s accumulated return reaches approximately $68$, which is $1.41\times$ that of RSPAR, $2.6\times$ greater than FMOLD, $2.62\times$ better than R2LTO and RLBR, $2.83\times$ higher than SMORLR, and $3.1\times$ higher than RLBEEP. Consequently, regardless of the exploration policy or preference variation patterns, the proposed method outperforms the baseline algorithms in terms of PDR.
VII-C Simultaneous Exploration and Exploitation